OpenStudy (anonymous):

Let A is a 4x4 matrix such that A(0,1,0,1)^t=(0,1,0,1)^t and A(1,0,1,0)^t=(-1,0,-1,0)^t. Also, suppose A has trace 4 and det 9/4. Find all eigenvalues of A and their multiplicites and explain why (1,1,1,1) is not an eigenvector of A.

OpenStudy (phi):

the two equations A(0,1,0,1)^t=(0,1,0,1)^t and A(1,0,1,0)^t=(-1,0,-1,0)^t. tell you 2 of the eigenvalues (and vectors): 1 and -1 the trace is the sum of the eigenvalues the set is the product of the eigenvalues

OpenStudy (phi):

**the determinant is the product of the eigenvalues

OpenStudy (anonymous):

All i'm confused about it the information at the beginning

OpenStudy (anonymous):

of the question

OpenStudy (anonymous):

how do you already know two eigen values?

OpenStudy (phi):

A(0,1,0,1)^t=(0,1,0,1)^t where t means transpose is a way to write \[A \left[\begin{matrix}0 \\ 1\\0 \\1\end{matrix}\right]= 1 \cdot \left[\begin{matrix}0 \\ 1\\0 \\1\end{matrix}\right]\]

OpenStudy (anonymous):

yea i understood that before.

OpenStudy (phi):

which should look familiar: \[ A v = \lambda v\]

OpenStudy (anonymous):

ohhhhhh

OpenStudy (anonymous):

yep. so the other would be -1. get ya

OpenStudy (anonymous):

yep now its easy to get lamda3 and lamda4

OpenStudy (phi):

yes. for the last part, [1 1 1 1] is a linear combination of the two eigenvectors they give you.

OpenStudy (anonymous):

yea. it looked obvious by looking at it from the beginning. thanks for starting me off!

OpenStudy (anonymous):

alright. so i found that lamda=1,-1,9/2,-1/2

OpenStudy (anonymous):

then does this mean their multiplicities are just 1?

OpenStudy (phi):

yes, all eigenvalues are distinct, so you have 4 eigenvalues, with multiplicity of 1 for each (a repeated value would give a multiplicity > 1)

OpenStudy (anonymous):

yea. how would you represent in mathematical terms? or is what you wrote sufficient?

OpenStudy (phi):

what I wrote should be sufficient

OpenStudy (anonymous):

so with b) (1,1,1,1)=(0,1,0,1)+(1,0,1,0) hence this is a linear combination of the two basis hence not an eigen vector?

OpenStudy (phi):

yes

OpenStudy (anonymous):

is it right to say that (0,1,0,1) and (1,0,1,0) are eigenbasis's?

OpenStudy (phi):

yes

OpenStudy (anonymous):

but each of them are eigen vectors?

OpenStudy (phi):

yes, the 4 distinct eigenvectors (because you have 4 distinct eigenvalues) make up a basis in 4d

OpenStudy (anonymous):

legend.

OpenStudy (anonymous):

thanks phi

OpenStudy (loser66):

thanks phi